OpenStudy (anonymous):

related rates problem. Air is being pumped into a spherical balloon at a rate of 5cm^3/min. Determine the rate at which the radius of the balloon is increasing when the diameter of the balloon is 20cm.

OpenStudy (anonymous):

do you have the formula for the area of a sphere given the radius?

OpenStudy (anonymous):

yep. V= 4/3pi(r^3)

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

ok then take the derivative with respect to time and you get \[V'=4\pi r^2r'\]

OpenStudy (anonymous):

you are told that \[V'=5\] so solve for \[r'\]

OpenStudy (anonymous):

where did the 3 from the 4/3 go?

OpenStudy (anonymous):

\[\frac{d}{dr}r^3=3r^2\] so the 3 canceled

OpenStudy (anonymous):

so the 3 on 3r\[^{2}\] was canceled?

OpenStudy (anonymous):

chain rule says \[\frac{dV}{dt}=\frac{dV}{dr}\frac{dr}{dt}\]

OpenStudy (anonymous):

you know \[\frac{dV}{dt}=5\] because that is what you were told in the problem

OpenStudy (anonymous):

\[V=\frac{4\pi}{3}r^3\] so \[\frac{dV}{dr}=4\pi r^2\]

OpenStudy (anonymous):

ok. that sort of makes sense.

OpenStudy (anonymous):

and therefore you get \[\frac{dV}{dt}=\frac{dV}{dr}\frac{dr}{dt}\] so \[5=4\pi r^2 \frac{dr}{dt}\]

myininaya (myininaya):

you want to find r' when r=10

OpenStudy (anonymous):

you want \[\frac{dr}{dt}\] so replace r by 10 and solve for it

OpenStudy (anonymous):

ok. one sec..

OpenStudy (anonymous):

you get \[5=40\pi \frac{dr}{dt}\] etc

myininaya (myininaya):

?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

wait.. isn't r^2 100?

OpenStudy (anonymous):

maybe it should be \[5=400\pi r'\]

myininaya (myininaya):

\[5=400 \pi r'\]

OpenStudy (anonymous):

that makes more sense. :)

OpenStudy (anonymous):

what myininaya said. good thing she is here to catch my mistakes....

myininaya (myininaya):

thats what i'm good for

OpenStudy (anonymous):

(and vice versa!)

myininaya (myininaya):

yes

OpenStudy (anonymous):

now back to work for me!

OpenStudy (anonymous):

so now that I have 5= 400pi(dr/dt), what next?

myininaya (myininaya):

lol solve for r'

myininaya (myininaya):

divide both sides by 400pi

OpenStudy (anonymous):

so.. divide both sides by 400\[\pi\] to get: 5/400pi = dr/dt

myininaya (myininaya):

5/(400pi) you can reduced the fraction